Use the elimination method to solve the system of equation that has four variables.
w+x-y+z=0
w-2x-2y-z=-5
w-3x-y+z=4
2w-x-y+3z=7
w=?
x=?
y=?
z=?

Answer by Alan3354(69443)

Use the elimination method to solve the system of equation that has four variables.
w+x-y+z=0
w-2x-2y-z=-5
w-3x-y+z=4
2w-x-y+3z=7
==============
Eliminated the z terms 1st.
Add eqn 1 & 2
Subtract eqn 1 & 3
Multiply eqn 1 by 3, then subtract from eqn 4.
----------
z is eliminated. You'll have 3 eqns in w, x & y.
Pick w, x or y and eliminate one of them.
etc.